OpenStudy (anonymous):

4 times the measure of the complement of an angle is equal to 2/3 the measure of the original angle's supplement. What is the measure of the angle?

OpenStudy (anonymous):

\[4\times(90°-x)=\frac{ 2 }{ 3 }(180°-x)\]

OpenStudy (muzzack):

wel if you want to solve for x isolate x in one side

OpenStudy (anonymous):

360-4x=120-2/3x 360-120=-2/3x+4x 240=10/3x / *3 720=10x /10 x=72 So it's 72 degrees.